# Diffusion Models for Robotic Manipulation: A Survey

### v1

@@ 2.2.1 Improving Sampling Speed and Quality
One group of methods aimed at improving sampling speed (Jolicoeur-Martineau et al.,, 2021; Song et al., 2021a, ; Lu et al.,, 2022; Karras et al.,, 2022) designs samplers that operate independently of the specific training process. Using an SDE/ODE-based formulation, allows choosing different discretizations of the reverse process than for the forward process. Larger step sizes reduce computational cost and sampling time but introduce greater truncation error. The sampler operates independently of the specific noise prediction network implementation, enabling the use of a single network, such as one trained with DDPM, with different samplers.
Denoising Diffusion Implicit Models (DDIM) (Nichol and Dhariwal,, 2021) is the dominant method used for robotic manipulation. It uses a deterministic sampling process and outperforms DDPM when using only a few (10-100) sampling iterations. DDIM can be formulated as a first-order ODE solver. In Diffusion Probabilistic Models-solver (DPM-solver) (Lu et al.,, 2022), a second-order ODE solver is applied, which decreases the truncation error, thus further increasing performance on several image classification benchmarks for a low number of sampling steps. In contrast to DDIM, Karras et al., (2022); Lu et al., (2022) use non-uniform step sizes in the solver. In a detailed analysis Karras et al., (2022) empirically show, that compared to uniform step-sizes, linear decreasing step sizes during denoising lead to increased performance (Karras et al.,, 2022), indicating that errors near the true distribution have a larger impact.
@@
There are also several methods (Zhou et al., 2024a, ; Li et al., 2024c, ; Wang et al., 2023b, ; Chen et al.,, 2024) regarding sampling speed, specifically for applications in robotic manipulation, which is different from the previously named methodologies, which were developed in the context of image processing. For example, (Chen et al.,, 2024) samples from a more informed distribution than a Gaussian. They point out that even initial distributions approximated with simple heuristics result in better sample quality, especially when using few diffusion steps or when only a limited amount of data is available.
@@ 2.3 Adaptations for Robotic Manipulation
Secondly, unlike in image generation, where the pixels are spatially correlated, in trajectory generation for robotic manipulation, the samples of a trajectory are temporally correlated. On the one hand, generating complete trajectories may not only lead to high inaccuracies and error accumulation of the long-horizon predictions, but also prevent the model from reacting to changes in the environment. On the other hand, predicting the trajectory one action at a time increases the compounding error effect and may lead to frequent switches between modes. Accordingly, trajectories are mostly predicted in subsequences, with a receding horizon, e.g., (Chi et al.,, 2023; Scheikl et al.,, 2024), which will be discussed in more detail in section 4.1.
@@ 3.1.1 Convolutional neural networks
The most frequently employed architecture is the CNN, more specifically the Temporal U-Net that was first introduced by Janner et al., (2022) in their algorithm Diffuser, a DM for robotics tasks. The U-Net architecture (Ronneberger et al.,, 2015) has shown great success in image generation with DMs, e.g., (Ho et al.,, 2020; Dhariwal and Nichol,, 2021; Song et al., 2021b, ). Thus, it has been adapted to robotic manipulation by replacing two-dimensional spatial convolutions with one-dimensional temporal convolutions (Janner et al.,, 2022).
The temporal U-Net is further adapted by Chi et al., (2023) in their CNN-based Diffusion Policy (DP) for robotic manipulation. While in Diffuser the state and action trajectories are jointly denoised, only the action trajectories are generated in DP. To ensure temporal consistency, the diffusion process is conditioned on a history of observations using feature-wise linear modification (FiLM) (Perez et al.,, 2018). This formulation allows for an extension to different and multiple conditions by concatenating them in feature space before applying FiLM (Li et al., 2024c, ; Si et al.,, 2024; Ze et al.,, 2024; Li et al.,, 2025; Wang et al., 2024b, ). Moreover, it also enables the incorporation of constraints embedded with an MLP (Ajay et al.,, 2023; Zhou et al.,, 2023; Power et al.,, 2023).
Discussed in more detail in section 4.1.1, Janner et al., (2022) formulates conditioning as inpainting, where during inferences at each denoising step, specific states from the currently being generated sample are replaced with states from the condition. For example, the final state of a generated trajectory may be replaced by the goal state, for goal-conditioning. This only affects the sampling process at inference and, thus, does not require any adaptations of the network architecture. However, it only supports point-wise conditions, severely limiting its applications. Multiple frameworks (Saha et al.,, 2024; Carvalho et al.,, 2023; Wang et al., 2023b, ; Ma et al., 2024b, ) directly employ the temporal U-Net architecture introduced by Janner et al., (2022). However, as this type of conditioning is highly limited in its applications, FiLM conditioning is more common. A different but less used architecture incorporates conditions via cross-attention mapped to the intermediate layers of the U-Net (Zhang et al., 2024a, ), which is more complicated to integrate than FiLM conditioning.
@@ 3.1.2 Transformers
Another commonly used architecture for the denoising network are transformers. A history of observations, the current denoising time step, and the (partially denoised) action are input tokens to the transformer. Additional conditions can be integrated via self-and cross-attention e.g., (Chi et al.,, 2023; Mishra and Chen,, 2024). The exact architecture of the transformer varies across methods. The more commonly used model is a multi-head cross-attention transformer as the denoising network , e.g., (Chi et al.,, 2023; Pearce et al.,, 2022; Wang et al., 2023b, ; Mishra and Chen,, 2024). Others (Bharadhwaj et al., 2024b, ; Mishra et al.,, 2023) use architectures based on the method Diffusion Transformers (Peebles and Xie,, 2023), which is the first method combining DMs with transformer architectures. There are also less commonly used architectures, such as using the output tokens of the transformer as input to a MLP, which predicts the noise (Ke et al.,, 2024).
For completeness we provide a list of works, using transformer architectures: (Chi et al.,, 2023; Pearce et al.,, 2022; Scheikl et al.,, 2024; Wang et al., 2023b, ; Ze et al.,, 2024; Feng et al.,, 2024; Bharadhwaj et al., 2024b, ; Mishra et al.,, 2023; Liu et al., 2023b, ; Xu et al.,, 2024; Mishra and Chen,, 2024; Liu et al., 2023c, ; Vosylius et al.,, 2024; Reuss et al.,, 2023; Iioka et al.,, 2023; Huang et al., 2025b, ).
@@ 3.1.3 Multi-Layer Perceptrons
Predominantly for application in RL, MLPs are used as denoising networks, e.g., (Suh et al.,, 2023; Ding and Jin,, 2023; Pearce et al.,, 2022), that take the concatenated input features, such as the observations, actions, and denoising time steps, to predict the noise. Although the architectures vary, it is common to use a relatively small number of hidden layers (2-4) Wang et al., 2023a ; Kang et al., (2023); Suh et al., (2023); Mendez-Mendez et al., (2023), using e.g., Mish activation (Misra,, 2019), following the first method (Wang et al., 2023a, ), integrating DMs with Q-learning. It is important to note that most of these methods do not use visual input. An exception from this is (Pearce et al.,, 2022), which also evaluates using high-resolution image inputs with an MLP-based DM. However, for this, a CNN-based image encoder is first applied to the raw image observation, before the encoding is fed to the DM.
@@ 3.1.4 Comparison
In summary, transformers are the most powerful architecture for handling high-dimensional input and output spaces, followed by CNNs, while MLPs have the highest computational efficiency. For processing visual data, such as raw images, an important task in robotic manipulation, a CNN or Transformer architecture should be chosen. Also, while MLPs are most computationally efficient, real-time control is possible with the other two architectures, integrating, for example, receding horizon control (Mattingley et al.,, 2011) in combination with a more efficient sampling process, like DDIM.
@@ 4.1 Trajectory generation
Trajectory planning in robotic manipulation is vital for enabling robots to move from one point to another smoothly, safely, and efficiently while adhering to physical constraints, like speed and acceleration limits, as well as ensuring collision avoidance. Classical planning methods, like interpolation-based and sampling-based approaches, can have difficulty handling complex tasks or ensuring smooth paths. For instance, Rapidly Exploring Random Trees (Martinez et al.,, 2023) might generate trajectories with sudden changes because of the discretization process. As already discussed in the introduction, although popular data-driven approaches, such as GMMs and EBMs, theoretically pertain to the ability to model multi-model data distributions, in reality, they show suboptimal behavior, such as biasing modes or lack of temporal consistency (Chi et al.,, 2023). In addition, GMMs can struggle with high-dimensional input spaces (Ho et al.,, 2020). Increasing the number of components and covariances also increases the models’ ability to model more complex distributions and capture complex and intricate movement patterns. However, this can negatively impact the smoothness of the generated trajectories, making GMMs highly sensitive to their hyperparameters. In contrast, denoising DMs have demonstrated exceptional performance in processing and generating high-dimensional data. Furthermore, the distributions generated by denoising DMs are inherently smooth (Ho et al.,, 2020; Sohl-Dickstein et al.,, 2015; Chi et al.,, 2023). This makes DMs well-suited for complex, high-dimensional scenarios where flexibility and adaptability are required. While most methodologies that apply probabilistic DMs to robotic manipulation focus on imitation learning, they have also been adapted to their application in RL, e.g., (Janner et al.,, 2022; Wang et al., 2023a, ).
In the following sections, the methodologies of DMs for trajectory generation will be further discussed and categorized. We will first explain their applications in imitation learning, followed by a discussion on their use in reinforcement learning.
@@ 4.1.1 Imitation Learning
In imitation learning (Zare et al.,, 2024), robots attempt to learn a specified task by observing multiple expert demonstrations. This paradigm, commonly known as Learning from Demonstrations (LfD), involves the robot observing expert examples and attempting to replicate the demonstrated behaviors. In this domain, the robot is expected to generalize beyond the specific demonstrations, which allows the robot to adapt to variations in tasks or changes in configuration spaces. This may include diverse observation perspectives, altered environmental conditions, or even new tasks that share structural similarities with those previously demonstrated. Thus, the robot must learn a representation of the task that allows flexibility and skill acquisition beyond the specific scenarios it was trained on. Recent advancements in applying DMs to learn visuomotor policies (Chi et al.,, 2023) enable the generation of smooth action trajectories by modeling the task as a generative process conditioned on sensory observations. Diffusion-based models, initially popularized for high-dimensional data generation such as images and natural languages, have demonstrated significant potential in robotics by effectively learning complex action distributions and generating multi-modal behaviors conditioned on task-specific inputs. For instance, combining with recent progress in multiview transformers (Gervet et al.,, 2023; Goyal et al.,, 2023) that leverage the foundation model features (Radford et al.,, 2021; Oquab et al.,, 2023), 3D diffuser actor (Ke et al.,, 2024) integrates multi-modal representations to generate the end-effector trajectories. As another example, GNFactor (Ze et al.,, 2023) renders multiview features from Stable Diffusion (Rombach et al., 2022b, ) to enhance 3d volumetirc feature learning.
In terms of the type of robotic embodiment, most works use parallel grippers or simpler end-effectors. However, few methods perform dexterous manipulation using DMs (Si et al.,, 2024; Ma et al., 2024a, ; Ze et al.,, 2024; Chen et al.,, 2024; Wang et al., 2024a, ; Freiberg et al.,, 2025), to facilitate their stability and robustness, also in this high-dimensional setting.
In the following sections, we will first repeat the process of sampling actions for trajectory planning with DMs and discuss common pose representations. Then we shortly address different visual data modalities, in particular 2D vs 3D visual observations. Afterwards, we look at methods formulating trajectory planning as image generation, before looking at applications in hierarchical, multi-task, and constrained planning, also looking at multi-task planning with vision language action models (VLAs). A visualization of the taxonomy is provided in table 1.
@@
As briefly discussed in section 2.3, the entire trajectory can be generated as a single sample, multiple subsequences can be sampled using receding horizon control, or the trajectory can be generated by sampling individual steps. Only in a few methods (Janner et al.,, 2022; Ke et al.,, 2024) the whole trajectory is predicted at once. Although this enables a more efficient prediction, as the denoising has to be performed only once, it prohibits adapting to changes in the environment, requiring better foresight and making it unsuitable for more complex task settings with dynamic or open environments. On the other hand, sampling of individual steps increases the compounding error effect and can negatively affect temporal correlation. Instead of predicting micro-actions, some use DMs to predict waypoints (Shi et al.,, 2023). This can decrease the combounding error, by reducing the temporal horizon. However, it relies on preprocessing or task settings, that ensure that the space in between waypoints is not occluded. Thus, typically, DMs generate trajectories consisting of sequences of micro-actions represented as end-effector positions, generally encompassing translation and rotation depending on end-effector actuation (Chi et al.,, 2023; Ze et al.,, 2024; Xu et al.,, 2023; Li et al., 2024c, ; Si et al.,, 2024; Scheikl et al.,, 2024; Ke et al.,, 2024; Ha et al.,, 2023). Although more commonly applied in grasp prediction, here the pose is sometimes also represented in special Euclidean group ($\mathbf{SE}(3)$) (Xian et al.,, 2023; Liu et al., 2023c, ; Ryu et al.,, 2024). Explained in more detail in section 4.2, the group structure of the $\mathbf{SE}(3)$ Lie group enables continuous interpolation and transformations between multiple object poses. As (Liu et al., 2023c, ; Ryu et al.,, 2024) performs complex tasks involving trajectory planning and grasping for aligning multiple objects, these properties are important to ensure physically and geometrically grounded actions. However, as the prediction of $\mathbf{SE}(3)$ poses with DMs requires a more complex model structure and training in imitation learning, it is more usual to use representations, such as euler angles or quaternions in trajectory planning. Once the trajectory is sampled, the proximity of the predicted positions enables computing the motion between the positions with simple positional controllers without the need for complex trajectory planning techniques.
@@
As already discussed in section 2.3 to ground the robots actions in the physical world, they are dependent on sensory input. Here, in the majority of methods visual observations are used. While in the original work (Chi et al.,, 2023), combining visual robotic manipulation with DMs for trajectory planning, RGB-images are used, this does not provide sufficient geometrical information for intricate robotic tasks, especially in scenes containing occlusions. Thus, multiple later methods used 3D scene representations, instead. Here, DMs are either directly conditioned on the point cloud (Li et al.,, 2025; Liu et al., 2023c, ; Wang et al., 2024a, ) or on point cloud feature embeddings (Ze et al.,, 2024; Xian et al.,, 2023; Ke et al.,, 2024), from singleview (Ze et al.,, 2024; Li et al.,, 2025; Wang et al., 2024a, ), or multiview camera setups (Ke et al.,, 2024; Xian et al.,, 2023). While multiview camera setups provide more complete scene information, they also require a more involved setup, and more hardware resources.
These models outperform methods relying solely on 2D visual information, on more complex tasks, also demonstrating robustness to adversarial lightning conditions.

### v2

@@ 2.2.1 Improving Sampling Speed and Quality
One group of methods aimed at improving sampling speed (Jolicoeur-Martineau et al.,, 2021; Song et al., 2021a, ; Lu et al.,, 2022; Karras et al.,, 2022) designs samplers that operate independently of the specific training process. Using an SDE/ODE-based formulation allows choosing different discretizations of the reverse process than for the forward process. Larger step sizes reduce computational cost and sampling time but introduce greater truncation error. The sampler operates independently of the specific noise prediction network implementation, enabling the use of a single network, such as one trained with DDPM, with different samplers.
Denoising Diffusion Implicit Models (DDIM) (Nichol and Dhariwal,, 2021) is the dominant method used for robotic manipulation. It uses a deterministic sampling process and outperforms DDPM when using only a few (10-100) sampling iterations. DDIM can be formulated as a first-order ODE solver. In Diffusion Probabilistic Models-solver (DPM-solver) (Lu et al.,, 2022), a second-order ODE solver is applied, which decreases the truncation error, thus further increasing performance on several image classification benchmarks for a low number of sampling steps. In contrast to DDIM, Karras et al., (2022); Lu et al., (2022) use non-uniform step sizes in the solver. In a detailed analysis Karras et al., (2022) empirically shows that compared to uniform step-sizes, linear decreasing step sizes during denoising lead to increased performance (Karras et al.,, 2022), indicating that errors near the true distribution have a larger impact.
@@
There are also several methods (Zhou et al., 2024a, ; Li et al., 2024c, ; Wang et al., 2023c, ; Chen et al.,, 2024) regarding sampling speed, specifically for applications in robotic manipulation, which is different from the previously named methodologies, which were developed in the context of image processing. For example, Chen et al., (2024) samples from a more informed distribution than a Gaussian. They point out that even initial distributions approximated with simple heuristics result in better sample quality, especially when using few diffusion steps or when only a limited amount of data is available. Others (Prasad et al.,, 2024) use teacher–student distillation techniques (Tarvainen and Valpola,, 2017), where pretrained diffusion models serve as teachers, guiding student models to operate with larger denoising steps while preserving consistency with the teacher’s results at smaller steps. While this increases training effort, it decreases sampling time at inference, which is especially important in (near) real-time control.
Recently, flow matching (Lipman et al.,, 2023) has been used as an alternative method to diffusion. Like with diffusion, the true distribution is estimated starting from a noise distribution. However, instead of learning the time-dependent score or noise, and then deriving the velocity from noise to data distribution from it, in flow matching, the time-dependent velocity field is learned directly. This leads to a simpler training objective, using the interpolation between the noise sample and true data point, without requiring a noise schedule. Thus, flow matching is usually more numerically stable and requires less hyperparameter tuning. However, when using few sampling steps, with flow matching, there is a risk of mode-collapse and infeasible solutions, as the ODE-solver averages over the velocity field. Thus, Frans et al., (2025) conditions the model not only on the time-step, but also on the step-size. By using the fact that one large step should lead to the same point as two consecutive steps of half the size, they maximize a self-consistency objective in addition to the flow-matching objective. Thus, the model can sample with a single step, with only a small drop in performance, far surpassing the performance of DDIM, when only a small number of sampling steps are used. While this is similar to the above-mentioned distillation techniques (Prasad et al.,, 2024), here only a single model has to be trained.
@@ 2.3 Adaptations for Robotic Manipulation
Secondly, unlike in image generation, where the pixels are spatially correlated, in trajectory generation for robotic manipulation, the samples of a trajectory are temporally correlated. On the one hand, generating complete trajectories may not only lead to high inaccuracies and error accumulation of the long-horizon predictions, but also prevent the model from reacting to changes in the environment. On the other hand, predicting the trajectory one action at a time increases the compounding error effect and may lead to frequent switches between modes. Accordingly, trajectories are mostly predicted in subsequences, with a receding horizon, e.g., (Chi et al.,, 2023; Scheikl et al.,, 2024), which will be discussed in more detail in Section 4.1 and is visualized in Fig. 2. In receding horizon control, the diffusion model generates only a subtrajectory with each backward pass. The subtrajectory is executed before generating the next subtrajectory on the updated observations. In comparison, grasps are generated similarly to images. As here only a single action, usually the grasp pose, is generated, this is done using a single backward pass of the diffusion model. Moreover, the grasp pose is usually predicted from a single initial observation. During execution, possible changes in the scene are not being taken into account. The backward pass for generating one action is visualized in Fig. 1.
Figure 2: Illustrations of the iterative trajectory generation using receding horizon control. At inference, the trajectory is planned up to a planning horizon $H$, conditioned on the past $P$ observations $\{o_{t},o_{t-1},\cdots,o_{t-P}\}$. Of this plan, only the steps until the control horizon $H_{c}\leq H$ are executed. In the figure, this is visualized in the outer loop with the time variable $t$. In the inner denoising loop, one subtrajectory $\tau=\{\tau_{t},\tau_{t+1},\cdots,\tau_{t+H}\}$ at the current time step $t$ is generated, using a diffusion model. Conditioned on the last $P$ observations and the current noise level $k$, the diffusion model predicts the noise, or score, dependent on the model type. Using the predicted noise/score, the trajectory at the next lower noise level $k-1$ is calculated. This is then used as the next input to the diffusion model until the trajectory is completely denoised ($k=0$), at which point it is executed. After execution of the subtrajectory, the time is increased and the next $H$ steps of the trajectory are planned. For training, ground truth trajectories and corresponding observations are sampled from the data buffer. The diffusion model is also trained on subtrajectories. However, the lookahead $H$ during training may be chosen larger than during inference, to ensure flexibility. The diffusion model is trained to predict the noise of a noisy trajectory. For this, first, a noise level $k$ is sampled. Then the noise $\epsilon_{k}$ is sampled, according to the predefined variance schedule. The noise is added in closed form to the ground-truth trajectory $\tau_{0}$ (see Eq. 4) to get the noisy trajectory $\tau_{k}$. The predicted noise $\epsilon_{\theta}(\tau_{k},k)$ on the trajectory $\tau_{k}$ is compared with the true sampled noise $\epsilon_{k}$ to compute the loss. Using this, the diffusion model can be updated.
@@ 3.1.1 Convolutional neural networks
The most frequently employed architecture is the CNN, more specifically the Temporal U-Net that was first introduced by Janner et al., (2022) in their algorithm Diffuser, a DM for robotics tasks. The U-Net architecture (Ronneberger et al.,, 2015) has shown great success in image generation with DMs, e.g., (Ho et al.,, 2020; Dhariwal and Nichol,, 2021; Song et al., 2021b, ). U-net, in general, is proven to be sample efficient and can even generalize well with small training datasets (Meyer-Veit et al., 2022b, ; Meyer-Veit et al., 2022a, ). Thus, it has been adapted to robotic manipulation by replacing two-dimensional spatial convolutions with one-dimensional temporal convolutions (Janner et al.,, 2022).
The temporal U-Net is further adapted by Chi et al., (2023) in their CNN-based Diffusion Policy (DP) for robotic manipulation. While in Diffuser, the state and action trajectories are jointly denoised, only the action trajectories are generated in DP. To ensure temporal consistency, the diffusion process is conditioned on a history of observations using feature-wise linear modification (FiLM) (Perez et al.,, 2018). This formulation allows for an extension to different and multiple conditions by concatenating them in feature space before applying FiLM (Li et al., 2024c, ; Si et al.,, 2024; Ze et al.,, 2024; Li et al.,, 2025; Wang et al., 2024b, ). Moreover, it also enables the incorporation of constraints embedded with an MLP (Ajay et al.,, 2023; Zhou et al.,, 2023; Power et al.,, 2023).
Discussed in more detail in Section 4.1.1, Janner et al., (2022) formulates conditioning as inpainting, where during inferences at each denoising step, specific states from the currently being generated sample are replaced with states from the condition. For example, the final state of a generated trajectory may be replaced by the goal state, for goal-conditioning. This only affects the sampling process at inference and, thus, does not require any adaptations of the network architecture. However, it only supports point-wise conditions, severely limiting its applications. Multiple frameworks (Saha et al.,, 2024; Carvalho et al.,, 2023; Wang et al., 2023c, ; Ma et al., 2024b, ) directly employ the temporal U-Net architecture introduced by Janner et al., (2022). However, as this type of conditioning is highly limited in its applications, FiLM conditioning is more common. A different but less-used architecture incorporates conditions via cross-attention mapped to the intermediate layers of the U-Net (Zhang et al., 2024a, ), which is more complicated to integrate than FiLM conditioning.
@@ 3.1.2 Transformers
Another commonly used architecture for the denoising network are transformers. A history of observations, the current denoising time step, and the (partially denoised) action are input tokens to the transformer. Additional conditions can be integrated via self-and cross-attention, e.g., (Chi et al.,, 2023; Mishra and Chen,, 2024). The exact architecture of the transformer varies across methods. The more commonly used model is a multi-head cross-attention transformer as the denoising network , e.g., (Chi et al.,, 2023; Pearce et al.,, 2022; Wang et al., 2023c, ; Mishra and Chen,, 2024). Others (Bharadhwaj et al., 2024b, ; Mishra et al.,, 2023) use architectures based on the method Diffusion Transformers (Peebles and Xie,, 2023), which is the first method combining DMs with transformer architectures. There are also less commonly used architectures, such as using the output tokens of the transformer as input to an MLP, which predicts the noise (Ke et al.,, 2024).
For completeness, we provide a list of works, using transformer architectures: (Chi et al.,, 2023; Pearce et al.,, 2022; Scheikl et al.,, 2024; Wang et al., 2023c, ; Ze et al.,, 2024; Feng et al.,, 2024; Bharadhwaj et al., 2024b, ; Mishra et al.,, 2023; Liu et al., 2023b, ; Xu et al.,, 2024; Mishra and Chen,, 2024; Liu et al., 2023c, ; Vosylius et al.,, 2024; Reuss et al.,, 2023; Iioka et al.,, 2023; Huang et al., 2025b, ).
@@ 3.1.3 Multi-Layer Perceptrons
Predominantly used for applications in RL, MLPs are employed as denoising networks, e.g., (Suh et al.,, 2023; Ding and Jin,, 2023; Pearce et al.,, 2022), which take concatenated input features, such as observations, actions, and denoising time steps, to predict the noise. Although the architectures vary, it is common to use a relatively small number of hidden layers (2-4) Wang et al., 2023b ; Kang et al., (2023); Suh et al., (2023); Mendez-Mendez et al., (2023), using e.g., Mish activation (Misra,, 2019), following the first method (Wang et al., 2023b, ), integrating DMs with Q-learning. It is important to note that most of these methods do not use visual input. An exception from this is (Pearce et al.,, 2022), which also evaluates using high-resolution image inputs with an MLP-based DM. However, for this, a CNN-based image encoder is first applied to the raw image observation, before the encoding is fed to the DM.
@@ 3.1.4 Comparison
In summary, transformers are the most powerful architecture for handling high-dimensional input and output spaces, followed by CNNs, while MLPs have the highest computational efficiency. For processing visual data, such as raw images, an important task in robotic manipulation, a CNN or a Transformer architecture should be chosen. Also, while MLPs are most computationally efficient, real-time control is possible with the other two architectures, integrating, for example, receding horizon control (Mattingley et al.,, 2011) in combination with a more efficient sampling process, like DDIM.
@@ 4.1 Trajectory generation
Trajectory planning in robotic manipulation is vital for enabling robots to move from one point to another smoothly, safely, and efficiently while adhering to physical constraints, like speed and acceleration limits, as well as ensuring collision avoidance. Classical planning methods, like interpolation-based and sampling-based approaches, can have difficulty handling complex tasks or ensuring smooth paths. For instance, Rapidly Exploring Random Trees (Martinez et al.,, 2023) might generate trajectories with sudden changes because of the discretization process. As already discussed in the introduction, although popular data-driven approaches, such as GMMs and EBMs, theoretically pertain to the ability to model multi-model data distributions, in reality, they show suboptimal behavior, such as biasing modes or lack of temporal consistency (Chi et al.,, 2023). In addition, GMMs can struggle with high-dimensional input spaces (Ho et al.,, 2020). Increasing the number of components and covariances also increases the models’ ability to model more complex distributions and capture complex and intricate movement patterns. However, this can negatively impact the smoothness of the generated trajectories, making GMMs highly sensitive to their hyperparameters. In contrast, denoising DMs have demonstrated exceptional performance in processing and generating high-dimensional data. Furthermore, the distributions generated by denoising DMs are inherently smooth (Ho et al.,, 2020; Sohl-Dickstein et al.,, 2015; Chi et al.,, 2023). This makes DMs well-suited for complex, high-dimensional scenarios where flexibility and adaptability are required. While most methodologies that apply probabilistic DMs to robotic manipulation focus on imitation learning, they have also been adapted to their application in RL, e.g., (Janner et al.,, 2022; Wang et al., 2023b, ).
In the following sections, the methodologies of DMs for trajectory generation will be further discussed and categorized. We will first explain their applications in imitation learning, followed by a discussion on their use in reinforcement learning. For an overview of the method architectures in imitation learning, see Table 2, and for reinforcement learning, see Table 3.
@@ 4.1.1 Imitation Learning
In imitation learning (Zare et al.,, 2024), robots attempt to learn a specified task by observing multiple expert demonstrations. This paradigm, commonly known as Learning from Demonstrations (LfD), involves the robot observing expert examples and attempting to replicate the demonstrated behaviors. In this domain, the robot is expected to generalize beyond the specific demonstrations, which allows the robot to adapt to variations in tasks or changes in configuration spaces. This may include diverse observation perspectives, altered environmental conditions, or even new tasks that share structural similarities with those previously demonstrated. Thus, the robot must learn a representation of the task that allows flexibility and skill acquisition beyond the specific scenarios it was trained on. Recent advancements in applying DMs to learn visuomotor policies (Chi et al.,, 2023) enable the generation of smooth action trajectories by modeling the task as a generative process conditioned on sensory observations. Diffusion-based models, initially popularized for high-dimensional data generation such as images and natural languages, have demonstrated significant potential in robotics by effectively learning complex action distributions and generating multi-modal behaviors conditioned on task-specific inputs. For instance, combining with recent progress in multiview transformers (Gervet et al.,, 2023; Goyal et al.,, 2023) that leverage the foundation model features (Radford et al.,, 2021; Oquab et al.,, 2023), 3D diffuser actor (Ke et al.,, 2024) integrates multi-modal representations to generate the end-effector trajectories. As another example, GNFactor (Ze et al.,, 2023) renders multiview features from Stable Diffusion (Rombach et al., 2022b, ) to enhance 3d volumetric feature learning. Very similar to diffusion, recently (Rouxel et al.,, 2024) flow-matching-based policies have emerged for trajectory generation, generally leading to a more stable training process with fewer hyperparameters, as already mentioned in Section 2.2.1. Nguyen et al., (2025) additionally includes second-order dynamics into the flow-matching objective, learning fields on acceleration and jerk to ensure smoothness of the generated trajectories.
In terms of the type of robotic embodiment, most works use parallel grippers or simpler end-effectors. However, few methods perform dexterous manipulation using DMs (Si et al.,, 2024; Ma et al., 2024a, ; Ze et al.,, 2024; Chen et al.,, 2024; Wang et al., 2024a, ; Freiberg et al.,, 2025; Welte and Rayyes,, 2025), to facilitate their stability and robustness, also in this high-dimensional setting.
In the following sections, we will first repeat the process of sampling actions for trajectory planning with DMs and discuss common pose representations. Then we shortly address different visual data modalities, in particular 2D vs 3D visual observations. Afterwards, we look at methods formulating trajectory planning as image generation, before looking at applications in hierarchical, multi-task, and constrained planning, also looking at multi-task planning with vision language action models (VLAs). A visualization of the taxonomy is provided in Table 1. More details on the individual method architectures are provided in Table 2.
@@
Reference Input Output Encoder Diffuser H
Chi et al., (2023) RGBMV RHC ResNet FiLM
Xian et al., (2023) RGB-D MV, Lan CT CLIP DiT & MLP
Reuss et al., (2023) GTS/RGB ${}^{\text{SV}}$ CT ResNet DiT
Chen et al., 2023a RGB ${}^{\text{SV}}$, Lan RHC ResNet U-Net
Zhou et al., (2023) RGB MV RHC CLIP U-Net
Pearce et al., (2022) RGB ${}^{\text{SV}}$ RHC CNN/ResNet MLP/DiT
Mendez-Mendez et al., (2023) GTS RHC - MLPs
Ze et al., (2024) PCs ${}^{\text{SV}}$ RHC MLP FiLM
[…]
Power et al., (2023) GTS RHC MLP U-Net
Ma et al., 2024b RGB-D ${}^{\text{SV}}$, Lan J PointNet++, MLP U-Net
Vosylius et al., (2024) RGB MV RHC Transformer DiT
[…]
Reuss et al., (2024) RGB MV, Lan RHC ResNet, CLIP DiT
Scheikl et al., (2024) RGB ${}^{\text{SV}}$/GTS RHC ResNet DiT
[…]
Li et al., 2024c RGBMV RHC ResNet FiLM
[…]
Saha et al., (2024) GTS RHC - U-Net
[…]
Wang et al., 2024b RGB, Tactile, PCs, Lan RHC ResNet, PointNet, T5 U-Net
Table 2: Technical details of trajectory diffusion using imitation learning. The references for the encoders are provided in Table 8.
In the following, the symbols and abbreviations are explained: H: Whether the method is hierarchical (✓) or not (✗). PCs: Point Clouds, Lan: Language, GTS: Ground Truth State, and wether the visual input modality is from single view (${}^{\text{SV}}$) or multi-view (${}^{\text{MV}}$). U-Net: temporal U-Net (Janner et al.,, 2022), FiLM: Convolutional Neural Networks with Feature-wise Linear Modulation Perez et al., (2018), DiT: Diffusion Transformer, RHC: sub-trajectories with receding horizon control, CT: complete trajectory in task space, J: complete trajectory in joint space. A “/” indicates that the information is not provided by the cited paper, while a “-” indicates that no specialized encoder is required as ground truth state information is used.
@@
As briefly discussed in Section 2.3, the entire trajectory can be generated as a single sample, multiple subsequences can be sampled using receding horizon control, or the trajectory can be generated by sampling individual steps. Only in a few methods (Janner et al.,, 2022; Ke et al.,, 2024) the whole trajectory is predicted at once. Although this enables a more efficient prediction, as the denoising has to be performed only once, it prohibits adapting to changes in the environment, requiring better foresight and making it unsuitable for more complex task settings with dynamic or open environments. On the other hand, sampling of individual steps increases the compounding error effect and can negatively affect temporal correlation. Instead of predicting micro-actions, some use DMs to predict waypoints (Shi et al.,, 2023). This can decrease the compounding error, by reducing the temporal horizon. However, it relies on preprocessing or task settings that ensure that the space in between waypoints is not occluded. Thus, typically, DMs generate trajectories consisting of sequences of micro-actions represented as end-effector positions, generally encompassing translation and rotation depending on end-effector actuation (Chi et al.,, 2023; Ze et al.,, 2024; Xu et al.,, 2023; Li et al., 2024c, ; Si et al.,, 2024; Scheikl et al.,, 2024; Ke et al.,, 2024; Ha et al.,, 2023). The control scheme is visualized in detail in Fig. 2. Although more commonly applied in grasp prediction, here the pose is sometimes also represented in special Euclidean group ($\mathbf{SE}(3)$) (Xian et al.,, 2023; Liu et al., 2023c, ; Ryu et al.,, 2024). Explained in more detail in Section 4.2, the group structure of the $\mathbf{SE}(3)$ Lie group enables continuous interpolation and transformations between multiple object poses. As (Liu et al., 2023c, ; Ryu et al.,, 2024) performs complex tasks involving trajectory planning and grasping for aligning multiple objects, these properties are important to ensure physically and geometrically grounded actions. However, as the prediction of $\mathbf{SE}(3)$ poses with DMs requires a more complex model structure and training in imitation learning, it is more usual to use representations, such as Euler angles or quaternions, in trajectory planning. Once the trajectory is sampled, the proximity of the predicted positions enables computing the motion between the positions with simple positional controllers without the need for complex trajectory planning techniques.
@@
As already discussed in Section 2.3 to ground the robots actions in the physical world, they are dependent on sensory input. Here, in the majority of methods, visual observations are used. While in the original work (Chi et al.,, 2023), combining visual robotic manipulation with DMs for trajectory planning, RGB-images are used, this does not provide sufficient geometrical information for intricate robotic tasks, especially in scenes containing occlusions. Thus, multiple later methods used 3D scene representations instead. Here, DMs are either directly conditioned on the point cloud (Li et al.,, 2025; Liu et al., 2023c, ; Wang et al., 2024a, ) or point cloud feature embeddings (Ze et al.,, 2024; Xian et al.,, 2023; Ke et al.,, 2024), from singleview (Ze et al.,, 2024; Li et al.,, 2025; Wang et al., 2024a, ), or multiview camera setups (Ke et al.,, 2024; Xian et al.,, 2023). While multiview camera setups provide more complete scene information, they also require a more involved setup and more hardware resources.
These models outperform methods relying solely on 2D visual information, on more complex tasks, also demonstrating robustness to adversarial lighting conditions.